OpenStudy (anonymous):

Where at on a number line does this go | x -1 | = 5

Parth (parthkohli):

\[|x - 1| = 5\]This means that \(x - 1\) is either \(5\), or \(-5\).

Parth (parthkohli):

If \(x - 1 = 5\), then \(x = 5 + 1 = 6\). What about \(x - 1 = -5\)?

OpenStudy (anonymous):

I would recommend: http://www.wolframalpha.com/input/?i= |+x+-1+|+%3D+5

OpenStudy (anonymous):

You do x-1 = 5 and x-1 = -5 and plot the answers to those on the number line and make your arrows.

OpenStudy (anonymous):

x-1=5 x=6 -x+1=5 -x=4 x=-4